Bando Corp. sells a single part for a price of $46 per unit. The variable costs of the part are $19 per unit and monthly fixed costs are $143,100.

Required:

a. What is the break-even level of monthly sales for Bando?

b. The cost analyst tells you that based on the price and cost information of the part and the marketing department's sales projection for next month, the margin of safety percentage is 20 percent. How many units does marketing expect to sell next month?
Answer #1

1. Break-even sales = Fixed cost/Contribution margin per unit

= 143,100/(46-19)

= 5,300 units

2. As we know that

Margin of Safety = Sales - Breakeven sales

= 5,300/(100-20)%

= 5,300/80%

= 6625 units